OpenStudy (anonymous):

A car traveled from town a to town b at the average speed of 24 miles per hour. the car then returned from town B to town A at the average speed of 48 miles per hour. What was the average speed of the car, on miles per hour, for the whole trip?

OpenStudy (anonymous):

one thing for sure, it is not \(\frac{24+48}{2}\)

OpenStudy (cwrw238):

its surprising how many people think that

OpenStudy (anonymous):

Thn

OpenStudy (anonymous):

lets call the distance between town A and B \(m\) for miles, there will be no \(m\) in the answer since distance is rate times time, we know the time going was \(\frac{m}{24}\) and the time returning was \(\frac{m}{48}\) therefore the total time was \[\frac{m}{24}+\frac{m}{48}=\frac{2m}{48}+\frac{m}{48}=\frac{3m}{48}\]

OpenStudy (anonymous):

the total distance back and forth was \(2m\) (since you went \(m\) mile going, and \(m\) miles returning

OpenStudy (anonymous):

so the average speed is total distance divided by total time, i.e. \[\frac{2m}{\frac{3m}{48}}\] invert and multiply, cancel whatever you can (including the \(m\) )and you will have your answer

OpenStudy (anonymous):

Please show me

OpenStudy (anonymous):

\[2m\times \frac{48}{3m}\] is a start

OpenStudy (anonymous):

Is it 72 by anychance

OpenStudy (anonymous):

no

OpenStudy (anonymous):

32

OpenStudy (anonymous):

the \(m\) cancels, so you get \[\frac{2\times 48}{3}\] now divide 3 in to 48, multiply the result by 2

OpenStudy (anonymous):

yes